An alpha particle, which contains two protons and
Answer #1


In moving throgh potential difference


work done done = change in KE

-delta_V*q = (1/2)*m*v^2

1000*2*1.6*10^-19 = (1/2)*4*1.6*10^-27*v^2

v = 3.16*10^5 m/s


=================


(b)

B = 1*10^-3 -k

v = 3.16*10^5 i m/s


Fb = q*(vxB)


Fb = 2*1.6*10^-19*( 3.16*10^5 i x 1*10^-3 -k )

Fb = 2*1.6*10^-19*3.16*10^5*1*10^-3 -(ixk)


Fb = 1.011*10^-16 N j

=======================

(c)


In the magnetic field


Fb = Fc


Fc = mv^2/r

1.011*10^-16 = 4*1.67*10^-27*(3.16*10^5)^2/r


r = 6.6 m

distance = pi*r = pi*6.6 = 20.73 m

=============

(d)

t = distance/v = 20.73/(3.16*10^5) = 6.56*10^-5 s


=============

(e)


v = - 3.16*10^5 i

==============


f)


Fb is in down wards


so electric force Fe is upwards


for positive charge electric force is in the direction of electric field


Electric field is in upwards

==================
g)

Fe = E*q

E = electric field


Fb = Fe

1.011*10^-16 = E*2*1.6*10^-19

E = 316 N/C

==================